OpenStudy (anonymous):

Having trouble with solving these equations. If somebody could help me step by step, that'd be great.

OpenStudy (anonymous):

\[\frac{ x }{ 2 }+\frac{ x }{ 3 }= \frac{ 5x+6 }{ }\]

OpenStudy (anonymous):

Over 6, my bad.

OpenStudy (anonymous):